OpenStudy (anonymous):

if P(A) is independent of P(BnC) is P(A) independent of P(B) and of P(C) when P(B) and P(C) are not necessarily independent?

OpenStudy (anonymous):

is this the exact question as written? it makes no sense, \[P(A)\] is a number, as is \(P(A\cap B)\) numbers are not sets, they are numbers and as such it makes so sense to ask if two numbers are "independent"

OpenStudy (anonymous):

Thats not the question asked but, I guess i mean A is independent of the intersection of B n C. Does that make A independent of B and A independent of C?

OpenStudy (anonymous):

Here is the problem I'm trying to solve

OpenStudy (anonymous):

\[P(A\cup B)=P(A)+P(B)-P(A\cap B)\]

OpenStudy (anonymous):

you have all the numbers on the right, plug them in

OpenStudy (anonymous):

I don't have P(A) or P(B) though

OpenStudy (anonymous):

ooh i see lets see if we can find them, give me a minute i bet a venn diagram might work well

OpenStudy (anonymous):

Yea this questions really hard I think

OpenStudy (anonymous):

I tried doing something like that but I couldn't get enough things to cancel out.

OpenStudy (anonymous):

one second it is totally doable we need to start in the center with \[P(O\cap E\cap F)\]

OpenStudy (anonymous):

since you are told \(O\) and \(E\cap F\) are independent, that means \[P(O\cap E\cap F)=P(O)\times P(E\cap F)=0.5\times .125=.0625\]

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

now i am stuck again damn

OpenStudy (anonymous):

oh maybe i have an idea now

OpenStudy (anonymous):

nicee

OpenStudy (anonymous):

this is taking a lot of paper, maybe i am missing something obvious

OpenStudy (anonymous):

thats why i was asking my original question, because I thought I was missing something like that

OpenStudy (anonymous):

damn i keep getting one more variable than i can get equations!!

OpenStudy (anonymous):

ok i am stuck

OpenStudy (anonymous):

I made the assumption i was asking about and got an answer i think is too low.

OpenStudy (anonymous):

nvm i made a mistake

OpenStudy (anonymous):

ok i fixed it and i got .675 as my answer which seems pretty legit

OpenStudy (anonymous):

question 26 is identical except for the numbers there does seem to be another assumption made that we did not use

OpenStudy (anonymous):

Thats what i used

OpenStudy (anonymous):

perhaps it was in the wording i took "the need for orthodontal work is indpendent for the need of filling and extraction" to mean \[P(O|E\cap F)=P(O)\] but they seem to have interpreted it differently as independent one at a time

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

yes i interpreted it that way as well. what im wonderig is if they are both true. ill let you know if i figure it out. g2g though thanks for your help :)